OpenStudy (anonymous):

y = -4* 2x is a reflection of y = -1/4* 2x. T/F?

OpenStudy (jdoe0001):

well, let's find out, let's get the inverse of "y" or f(x), that is, let's get \(\large f^{-1}(x)\)

OpenStudy (jdoe0001):

from y = -4* 2x swap the letters about, what do you get?

OpenStudy (anonymous):

True?

OpenStudy (anonymous):

hmmm.... nope, looks false to me. y = -4*2x ===> y = -8x swap the letters x = -8y solve for y now y = -x/8 that doesn't look the same as y=-1/4 * 2x which is actually y = -1/2x

OpenStudy (anonymous):

yes you're right @robz8